Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 336     Entry time: 08 Sep 2023, 06:22
Author: Erik Brambrink 
Group: HED 
Subject: Shift summary 
  • Improved the alignment procedure:
    • Using the pin to make sure of the spatial overlap
    • SAXS signal is used to make sure the pin is well centered on the X-ray beam path
    • YAG is only used in combination with ILM to fix with ~5 um precision the height of the beam (moving CRLs on laser focus) and the temporal overlap.
  • Recording of the different laser foc planned to be used during the experiment: no aperture, 40 mm, 25 mm, 15 mm diameter

Note:

  • Jet failed a couple of time
  • We lost the Sync system for 2 hours, got it back again for 30 min, then lost it again
ELOG V3.1.4-7c3fd00